OpenStudy (tainted):

Show work please √(x+2+4)=x

OpenStudy (anonymous):

Simplify. Square both sides.

OpenStudy (anonymous):

Simplify x+2+4 = x+6. You then need to square both sides to get rid of the square root on the left \[(\sqrt{x+6})^{2}=(x)\] Then put all the terms on one side and factor by factoring or using the quadratic equation to find x

OpenStudy (anonymous):

Sorry i meant \[(\sqrt{x+6})^{2}=(x)^2\]